### v1

# Characterising and tackling thermally induced zero-drift in displacement measuring interferometry using temperature-controlled enclosure

Abstract
Our research efforts in displacement measurement interferometry focused on long-term drifts initiated an extended experimental investigation in the interferometric assemblies of our design. We aimed to analyze, characterize and tackle the long-term measurement stability, expressed as the zero-drift, with special attention to the thermal effects. For the experimentation, we developed a thermostatic chamber equipped with an active temperature regulation, array of sensors and control electronics. With either the finely stabilized temperature or with the thermal cycling, we are able to carry out a range of investigations: verification of modified design or prototype interferometers, testing of production pieces, characterization of integrated assemblies and units in terms of the zero drift and the susceptibility to thermal effect - the temperature sensitivity, expressed as in nm.K.
Note: This is the version of the article before peer review or editing, as submitted by an author to Measurement Science and Technology. IOP Publishing Ltd is not responsible for any errors or omissions in this version of the manuscript or any version derived from it.
Keywords: displacement measuring interferometry, dimensional metrology, thermal compensation, zero-drift, thermal drift
1 Introduction
Interferometry, a cornerstone technique in modern length metrology, has revolutionized the field of dimensional measurement by offering unparalleled precision and accuracy[1]. Employing the principles of wave interference[2], interferometry enables the determination of distances and dimensions with extraordinary sensitivity, making it indispensable in a wide range of scientific and industrial applications. At the same time, interferometric methods have continually pushed the boundaries of achievable precision.
With the resolution and precision that matches and even dives below the size of single atoms, there is a naturally long list of aspects and influences that require careful consideration to achieve successful interferometric measurement[3]. Many of these aspects have already been well covered in the literature and are still actively investigated.
However, the question of the (long-term) stability of displacement reading, typically expressed as the zero-drift (e.g. in the ANSI/ASME B89.1.8 standard[4]), has been so far either overlooked (for many good practical reasons) or simply not of interest. Reducing the zero drift is generally desirable for any interferometric instrumentation. Nonetheless, there are two specific paths of development where minimal drifts are mainly of interest: measurement in an environment with sub-optimal conditions and measurement over extended time frames. In these areas, the zero-drift is an obvious obstacle to the broader penetration of interferometry to metrological applications in the high-tech industry and generally beyond the boundaries of national metrology laboratories.
The extended time-frame examples typically occur in applications that require precise and stable positioning (large area scans in nanometrology, long expositions in e-beam lithography, reference object positioning in optical metrology), in applications where the interferometer(s) serve as a length reference (large volume metrology, calibrations with larger than usual number of measurement points or repetitions) and also in fundamental research (e.g. research on gravitational waves [5], realization of the unit kilogram using Kibble balance [6]).
The other part seeks to deliver the precision, inherent traceability and dynamic range of interferometric measurements outside the laboratories with a well-controlled and monitored environment, primarily temperature, air pressure and relative humidity. The applications range from 100 characterization of sensors elements [7] in production (as required by Industry 4.0), via fast turnover in-house calibration of working standards (e.g. gauge blocks, end bars) to in-situ calibration of instrumentation for a highly specific task (such as structural inspection in nuclear power plants [8]).
Our work aims to investigate, reduce, and characterize the residuals of zero-drifts, especially thermally induced, to make the unprecedented accuracy of the interferometric measurement applicable to these kinds of applications and measurement scenarios.
In this paper, we explore the length drifts beginning with the perspective of the entire measurement system assembly, quickly moving towards the system’s heart, the interferometer itself. During the journey, we identify the main culprit: the omnipresent effects of thermal changes on both the geometrical and material characteristics (Section 2). With this in mind, we design and realize a dedicated temperature-controlled enclosure as an experimental testbed to investigate the interferometric components and assemblies under thermal load (Section 3). We present a variety of (currently ongoing) investigations and an example of the improvement in tackling the zero drift in interferometric systems (Section 4). The paper is then summarised and concluded with future prospects in what we see as a broad potential for progress in interferometry-based dimensional metrology (Section 5).
2 Thermal effects influencing interferometric measurement
Temperature and its fluctuations profoundly influence dimensional measurement and impact the measurements’ accuracy and reliability. This influence primarily stems from the thermal expansion, which alters the geometry of both the measuring arrangement (not necessarily interferometric at this point) and the measurand. While the expansion of the measurand is typically accounted for (e.g. in gauge-block calibration[9]), accounting for the dimensional changes that influence the structural loop and, consequently, the measurement loop of the measurement systems can be challenging and is often neglected. The most straightforward way to limit the thermal effects at the system level is to keep the outer environment stable (a remarkable approach could be found in Helsinki [10]) and let the measurement equipment stabilize well before the actual measurement. The disadvantage is that any local temperature disturbances (e.g. heat dissipation from electronics or translation mechanisms) interfere with the temperature control and induce uneven temperature distribution along the measurement assembly. Another (complementary rather than alternative) approach to mitigate the thermal issues incorporates careful design of the measuring system focused on a well-stabilized metrology loop so that the thermal creeps have limited influence on the measurement arrangement. Typical examples could be using low-expansion materials for the metrology frame [11], using differential interferometers [12, 13] to shorten the loop and, generally, sticking to good practice for precision engineering [14].
With interferometer(s) involved, the temperature has a massive influence on the refractive index of air. Unless the entire measurement system is put into vacuum [15], the measurement accuracy is hampered by K [16]. This influence is usually mitigated using a suitable empirical equation (i.e. indirect refractometry[17]). Nonetheless, the residual issues with a non-uniform temperature distribution of unknown magnitude could remain (for the comparison, the climate chambers used for such purposes usually declare the temperature homogeneity somewhere at C).
A further element of interest is the susceptibility to the thermal effects of the interferometer itself, particularly the interferometer’s optics. Of the two mentioned influences – the thermally induced geometry changes and changes in the refractive index – the interferometer is significantly subject to both. With a temperature variation, the geometrical dimensions change, and so do the refractive index of the optical material the interferometer’s optics is manufactured of. The variation of the refractive index with temperature is referred to as the thermo-optic coefficient [18, 19]. Table 1 presents selected optical materials and corresponding values of the coefficient of the thermal expansion (CTE) and the thermo-optic coefficient (dn/dT [19]). The values indicate that these influences have a magnitude comparable to that of the thermal expansion and the fluctuations of the refractive index of air. To minimize the thermal drifts, it is crucial to design the optical arrangement of the interferometer in a way the parts of the interferometer’s arm within the interferometer’s optics are of equal optical length[20] and ideally share the path to a maximal extent. Despite these findings being decades old, it seems many state-of-the-art interferometric designs tend to neglect this aspect [21, 13]. Also, the thermal drifts were not thoroughly experimentally characterized in detail at the time of original investigation.
CTE dn/dT Material (K) (K) Notes Schott N-BK7 7.1 1.29 Schott N-SF14 9.41 -1.46 Fused silica 0.55 12.9 Crystal quartz 7.5 -5.5 p-polarization 13.5 -6.5 s-polarization
Finally, the technique of optical / opto-mechanical assembly can also render the interferometer sensitive to temperature changes. An interferometer assembled from individual components (rather than cemented or contacted together [13]) could be potentially susceptible to thermal drifts due to uneven influence of local temperature or pressure change or some misalignments that would, e.g. not exhibit on the fringe contrast directly. This susceptibility will differ depending on the particular arrangement and vary piece by piece due to the manufacturing process.
In a real experiment or application, all the temperature-induced effect sum up and the resulting superpositions exhibits as the zero-drift. Due to non-ideal aspects of the reality, every compensation mechanism has its limits. It is inevitable that any interferometric measurement will suffer certain amount of residual thermal drift and exhibit some level of (dominantly thermally-induced) zero-drifts.
These residual drifts become more pronounced in the scenarios mentioned above with the measurement over a longer period of time and in an environment with sub-optimal thermal conditions. The challenge we accepted was to specify and investigate a methodology allowing us to experimentally test the components of interferometric systems in an environment with stable and controlled temperature. This approach is intended to characterize the residual thermal drifts and quantify them as a measure called the temperature sensitivity, expressed as a displacement depending on the temperature in .
3 Temperature-controlled experimental enclosure
Typical instrumentation used to achieve temperature control in an air environment is some kind of climate chamber. Besides the costs, the major disadvantage is the homogeneity of the temperature inside the box (usually ) in conjunction with active air circulation, which is generally not favourable for interferometric measurements.
3.1 Enclosure design
To overcome these limitations so that we can characterize the thermal sensitivity we have designed a temperature-controlled experimental enclosure with the goal to achieve thermal stability better than mK. This (intermittent) goal was chosen so that the relative accuracy of displacement measurement, burdened with the thermal expansion and fluctuations of the air refractive index, would be at a similar level as the other significant source of drifts: the frequency source ( with a frequency-stabilized He-Ne laser). The entire ecosystem, sketched in Figure 1, consists of the actual enclosure, set of environmental sensors and control electronics.
The walls of the enclosure with a volume of litres and are lined with a polyethene insulation foam to provide the passive insulation. The enclosure’s floor holds a breadboard (for versatile assembly of tested devices), and its honeycomb construction ensures mechanical stability during thermal loads. An array of thermo-electric cells (six pieces) are mounted on the enclosure ceiling. One side is thermally connected to passive coolers outside the box, and the other to a homogenization plate inside. The enclosure is designed to be operated with the temperature set slightly below the room temperature so that the peltier cells cool down the inner space from the top, relying on thermal convection.
The chamber is equipped with a range of sensors (see also Figure 3). Several temperature sensors monitor the temperature conditions. The servo temperature sensor (T) serves as a process variable for the servo-loop temperature control. The reference temperature (T) sensor with calibrated absolute scale monitors the temperature at the same point as the servo-temperature sensor. The other pair monitors te external temperature (T) outside the enclosure, and the gradient temperature (T) is placed above the monitor and servo sensors (vertically spaced mm ) to detect the vertical temperature gradient. Besides the temperature, the air’s barometric pressure and relative humidity are measured.
The control electronics comprise three integral modules. The temperature controller TC interrogates the servo temperature (with mK resolution and excellent relative stability) and drives the current to the thermo-electric cells (up to A). The interferometric modul IFM interrogates the quadrature sine-cosine output either from the homodyne optical receivers or similar interface for the case of commercial instrumentation tested in the chambers. The environmental sensor control unit E.I.U (originally an indirect refractometer[22]) handles the rest of the sensors: three temperature sensors with the resolution of mK were calibrated with K (range from C to C, ), the pressure sensor (resolution Pa, range from kPa to kPa, kPa) and sensor for the relative humidity (). The E.I.U is powered from separate module PWR. Due to precision limits, a separate reference barometer (Druck PACE 1000) with Pa precision was operated along the enclosure.
3.2 Thermal conditions in the enclosure
To obtain insight into the thermal conditions in the enclosure, we analyzed two weeks long recording from the temperature sensors, displayed in Figure 2. The temperature fluctuated with mK as measured by both the servo temperature sensor and the reference temperature sensor. The good relative coincidence of the two sensors (mK), which is at the level of resolution of the less precise sensor, proves there are no long-term drifts in the temperature measurement. The recording of the external temperature (outside the box) indicates that the external temperature and inner temperatures are coupled (and further tuning of the servo-control is necessary). Finally, the measurement revealed a negative vertical temperature gradient inside the enclosure of approximately K.m, which is quite significant and requires further investigation.
4 Experimental investigations with the enclosure
We present several use cases where the characterization of (residual) thermal drifts is proven to provide useful qualitative results. The standard arrangement of the experiments, shown in Figure 3, was (with only a slight variation) used across most of the individual experiments.
The servo-controlled temperature in the enclosure allowed for two types of investigation. The temperature is kept stabilized for a prolonged period (days to weeks) so that the limits of the thermo-mechanical stability of individual components or more complex assemblies can be investigated. The complementary approach (that proved more useful) was to apply the thermal load in predefined cycles with linear ramp-up. With the bulk interferometers, the feasible temperature change rate was identified as K / hour. This value corresponds to quite a small power rating of the cooling relative to the enclosure volume. A faster heating/cooling rate led to more destabilized thermal conditions in the enclosure and, in turn, distorted measurement data.
The following use cases represent several experimental scenarios that address different research questions revolving around studying the thermal drifts in length measurement instrumentation. Combined together, they give an indicative overview of the range of testing and characterization possibilities.
4.1 Comparing variants of interferometric arrangement
In this experiment, we have compared two two-beam double pass interferometers (non-differential; a standard style used e.g. for coordinate positioning). One of the interferometers had the optical arrangement slightly modified so that the optical paths of the individual interferometer’s arms within the bulk optics of the interferometru were of equal length (so that one interferometer had the beams’ optical paths of different lengths, and the other had them of equal length). This modification was expected to reduce the thermal drifts and the objective of the test was to confirm that the modification of the beam paths will have desired effect.
The interferometers were fixed on standard kinematic mounts (Thorlabs) and aligned against silver mirrors. The interferometers were loaded with a temperature change of K with the cooling/heating rate of K/hour. The observed drifts, shown in Figure 4, were highly correlated to the induced thermal cycling () and improved more than twofold reduction. Nonetheless, the residual drift of the ”balanced” interferometer still came out significant, and it could be primarily attributed to thermal expansion of the opto-mechanics used since the interferometer are not differential.
4.2 Identifying faulty optical assembly
The aim of the test was to investigate drifts of a differential interferometer (modified arrangement[13]) produced with a modified assembly procedure. The interferometer was tested in the enclosure (without the active thermal control at that time) to investigate the drifts using the approach shown in Figure 3.
The interferometer did not show any signs of malfunctions (such as poor contrast, increased noise, presence of etalon effects, or quadrature signal distortion). The repeated tests in the enclosure indicated significant drifts (see the upper part in Figure 5) of hundreds of nanometres (over several weeks) instead of expected fluctuation at a nanometre level (see the lower part in Figure 5). After all other components were gradually replaced to rule out their contribution, the interferometer underwent partial disassembly. A loosened single corner of a retarder waveplate was revealed and identified as the source of drifts.
The conclusion is that zero-drift testing has a significant potential not only in characterizing the drifts but also in verifying the conformance of the interferometers (and generally any measuring instruments). The prior knowledge that is obviously necessary could come either from the previous experimental results or from analytical sources (simulations, calculations). The two significant benefits are that the technique could reveal faults otherwise difficult to detect and that the approach is simplistic enough to be useful for 100% verification of optical systems.
4.3 Improving mounting stability
Another measurement scenario demonstrates that the range of testing possibilities with the enclosure is not limited to the interferometer optics. When asked to characterize a pair of commercial interferometers, we used the standard experimental arrangement (Figure 3) to test both simultaneously. The interferometers were mounted in a vertical orientation according to the mechanical interface specified in the device documentation. Generic opto-mechanical components (Thorlabs) were used as the baseplates.
With the rough idea of the differential double-pass interferometer’s arrangement, we expected fairly pronounced thermal drift. The test results, however, revealed a significant global drift of several hundred nanometres (see the upper part in Figure 6). In several consecutive attempts, the fixture was gradually replaced with more robust components with enhanced triangular supports and additional side supports were added. With these measures applied, the observed drift was reduced by a factor of five with one of the interferometers and even more with the second one of the pair (see the lower part in Figure 6). Notably, the effect of the thermal load was still less significant than the residual (global) drift.
With the analysis of thermal influences presented before in mind, this type of test could provide a useful tool in the characterization and verification of complex (sub-) systems intended to become part of measurement instrumentation. It is also evident that the individual manifestations of thermal changes are principally superimposed and challenging to discern, and their mitigation requires careful consideration in multiple iterations.
4.4 Comparison of the assembly technology
The last experimental investigation we present is similar to the comparison of interferometers with the slightly varied arrangements (Section 4.1). This time, we loaded the enclosure with three interferometers of identical arrangement but differing in the assembly technique and, consequently, the glass material of the bulk components. The first interferometer has the components separately glued into the metal (aluminium) housing (cage assembly). In the second case, the optical components were cemented together (glue assembly), and the last interferometer was assembled using optical contacting (wringing). The first two interferometers were made of the N-BK7 and N-SF14 glasses (Schott), while the last one was made of fused silica. Instead of the standard optomechanics, the interferometers were fixtures on a titanium baseplate with additional support for the mirror’s kinematic mount. Due to the improved fixture, the global drift has been significantly reduced. The results of the test, displayed in Figure 7, indicate the temperature sensitivity is between nm.K and nm.K. The interferometer assembled by optical contacting exhibits the highest sensitivity because of the material (remember Table 1): fused silica has a significantly higher thermo-optic coefficient, so the influence on the optical lengths is the most pronounced. The sensitivity of the other two interferometers is comparable. What remains unaccounted for is the thermal expansion of the titanium baseplate, so that the inferred temperature sensitivity figures shouldn’t be considered absolute.
5 Final words
We have presented the achieved status in designing and developing a thermal-controlled enclosure dedicated to testing and characterising displacement measuring interferometers and their assemblies. The servo-loop temperature control achieves temperature stability of during a day and enables setting the temperature in the range of several kelvins. The enclosure could be used to keep the temperature stable over days and weeks and also makes possible to apply controlled thermal load to the devices under test.
We demonstrated four use cases that used either of the capabilities for various characterization or testing tasks. The comparisons of zero-drift between similar interferometers (different assembly techniques, variant optical arrangement) provided valuable feedback for expanding the capabilities of and refining the interferometric measurement technology. Similarly, using the thermal load for intermittent testing provided feedback to redesign the interferometer’s mechanical fixtures, which significantly reduced the drifts. The enclosure’s functionality proved feasible for routine verification of the optical and optomechanical systems, where the testing under thermal load could help identify otherwise undetectable failures in the systems. Last but not least, the construction of the temperature-controlled enclosure and characterization of the thermal conditions inside could be transferred to the design of measurement instrumentation. The proper, intelligent and highly precise thermal management should be an intrinsic part of the design (e.g. [23]) not only in the top-grade instruments at national metrology laboratories – this an appealing approach to carrying out measurements under defined conditions, and simultaneously a necessary condition on the way to lower measurement uncertainty and an extended range of measurement scenarios.
While the particular results we presented tend to be rather indicative than absolute, the message and contribution of this paper to our ultimate belief are demonstrative and methodological. The current state of the work opens both apparent questions and challenges (e.g. to fine-tune the temperature control servo loop, to develop bases for non-differential interferometers from low-expansion material) as well as the broader questions with overlap to precision engineering and measurement technology (e.g. investigation of the temperature distribution in the thermally controlled environment, because the temperature stability still relies on a single point measurement). Back to the interferometry, we also believe this work is crucial on the way towards further expansion of interferometric technology to routine operation in longer timeframes or in a less controlled environment – for instance, in the heart of a nuclear power plant[8].
Acknowledgement
Technology Agency of the Czech Republic (FW03010687;TN02000020); European Commission (CZ.1.05/2.1.00/01.0017, LO1212, CZ.02.1.01/0.0/0.0/16_026/0008460); Czech Academy of Sciences (RVO:68081731)
Disclosure
The authors declare no conflicts of interest.
Data Availability Statement
The authors are willing to share the data that support the findings of this study on the basis of a reasonable request.
Authors’ Contribution
S. Rerucha– Conceptualization, Funding acquisition, Investigation, Methodology, Software, Data Curation, Validation, Visualization, Writing - Original Draft, Writing - Review and Editing. M. Hola – Investigation, Resources (optics, experimental setups). J. Oulehla – Resources (optical coatings), Investigation, Data Curation. J. Lazar – Conceptualization, Methodology. B. Mikel – Project management, Supervision. O. Cip – Conceptualization, Funding acquisition, Supervision. All authors – Writing - Review and Editing.